OpenStudy (anonymous):

Find the unknown side length for the rectangle. Area = 27 in² A. 3 in. B. 6 in. C. 9 in. D. 24 in. putting the pic below

OpenStudy (anonymous):



OpenStudy (anonymous):

@RCCB

OpenStudy (jango_in_dtown):

area=l*b here b=3 and area=27 so l=27/3=9

OpenStudy (anonymous):

@RCCB whats the answer

OpenStudy (anonymous):

@bibby

OpenStudy (jango_in_dtown):

9

OpenStudy (jango_in_dtown):

C. 9 in

OpenStudy (jango_in_dtown):